OpenStudy (anonymous):

Solve using the substitution method 3x-3y=-18 7x+72=y

OpenStudy (anonymous):

hey Rachel, yesterday I posted a detailed answer to one of your similar questions, I suggest you make an effort to learn to do these on your own.... You can do the steps and i will confirm if u hv done ok or not..... that will u will be able to solve them in a test also on your own..

OpenStudy (anonymous):

I apologize I am trying to get an assignment done. I live in a city in arizona that is under is in a state of emergency for a wild fire. I have been evacuated and this assignment is already past due. I am just trying to get the last few problems done that I am having problems with.

OpenStudy (anonymous):

its a system

OpenStudy (anonymous):

ok will help then..

OpenStudy (anonymous):

take y and subtract

OpenStudy (anonymous):

take the 72 to the other side by subtracting

OpenStudy (anonymous):

take the second equation and multiply by -3

OpenStudy (anonymous):

now add down

OpenStudy (anonymous):

3x+-21x =-18x

OpenStudy (anonymous):

the y's cancel

OpenStudy (anonymous):

-18+-72=-90

OpenStudy (anonymous):

reduce lst equation to x-y=-6 then we will get x=-6+y then we substitute it to second equation an we will get -42+7y+72=y,y=-5,x=-11

OpenStudy (anonymous):

-90/-18=10/3

OpenStudy (anonymous):

so x =10/3

OpenStudy (anonymous):

now you substitute 10 thirds into the equation to get your y coordinate

OpenStudy (anonymous):

given equations are 3x - 3y = -18 ------- (1) 7x - y = -72 ---------- (2) (this u get after rearranging) multiply (2) by 3, to get 21x - 3y = -216 now and subtract it from (1) to get - 18x = 198 x = 198/-18 = -11 now put x = -11 in (1) 3(-11) - 3y = -18 -33 - 3y = -18 -3y = -18 + 33 y = 15/-3 = -5

OpenStudy (anonymous):

@chris u hv made some errors, your values are wrong..... pls recheck

OpenStudy (anonymous):

I multiplied by -3 to cancel the y's

OpenStudy (anonymous):

the second shouldve been -21x +3y=

OpenStudy (anonymous):

216 my bad forgot to multiply the last number

OpenStudy (anonymous):

@chris pls check in yr ans u hv put -90/-18 as 10/3, this is incorrect -90/-18 = 5

OpenStudy (anonymous):

good lookin

OpenStudy (anonymous):

got a medal for me Rachel !!!!!

OpenStudy (anonymous):

harkerat help me with one

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

of course you are so thorough...didn't know it was you though since your name is different

OpenStudy (anonymous):

sorry chris, am busy with other things tooo... go on what can i help u with....